# LLMs as Bridges: Reformulating Grounded Multimodal Named Entity Recognition

### v3

@@ 5 Conclusion
In this paper, we propose the RiVEG, a unified framework that aims to use LLMs as a bridge to reformulate the GMNER task into a unified MNER-VE-VG task. This reformulation enables the framework to seamlessly incorporate the capabilities of state-of-the-art methods for each subtask in their corresponding modules. Extensive experiments show that RiVEG significantly outperforms state-of-the-art methods. The proposed RiVEG greatly reduces the constraints of VG and VE methods on natural language inputs, expands their applicability to more realistic scenarios, and provides insights for future work to better deploy VG and VE modules in related tasks.
@@ A.1 Model Configuration
A.2 Details of Different Module Datasets
| #Twitter-GMNER | #Twitter-GMNER | # | # | # | # | |
| Train | 7000 | 35000 | 11782 | 58910 | 4694 | 23470 |
| Dev | 1500 | 1500 | 2453 | 2453 | 986 | 986 |
| Test | 1500 | 1500 | 2543 | 2543 | 1036 | 1036 |
| Total | 10000 | 38000 | 16778 | 63906 | 6716 | 25492 |
Table 7 shows the details of the Twitter-GMNER dataset. And Table 8 counts the training data used by different modules. For the MNER module of RiVEG, the training data for base version comprises original samples and corresponding auxiliary refined knowledge generated by gpt-3.5-turbo provided by PGIM (Li et al., 2023). Additionally, data augmentation is applied to the training set of base version. In data augmentation version, we apply four additional LLMs (i.e., vicuna-7b-v1.5, vicuna-13b-v1.5, llama-2-7b-chat-hf, llama-2-13b-chat-hf) to training set samples using the same algorithm, generating multiple versions of auxiliary refined knowledge for the same sample. The validation set and test set remain the same as base version to ensure the benchmarking of evaluation.
@@ A.5 Exploration of More Visual Variants
We also present the corresponding test set results for different methods of different modules after fine-tuning. The results indicate that stronger combinations of sub-methods generally yield superior performance. The performance of the early classic ALBEF and SeqTR is inferior to the subsequent advanced OFA. Additionally, while the test set Acc@0.5 of SeqTR is slightly lower than that of OFA, there are significant differences in their final GMNER and EEG results. This phenomenon may be attributed to the text representation method employed by SeqTR. The GRU (Chung et al., 2014) vocabulary of SeqTR is solely statistically obtained from or , rendering it incapable of handling out-of-vocabulary words in new samples during inference.
Moreover, data augmentation methods exhibit varying effects on distinct modules. MNER and VG methods can significantly benefit from data augmentation, but this is not the case for VE methods. One potential reason is that the distribution after fitting more training data does not align with the optimal distribution of the test data. But this is acceptable since data augmentation significantly enhances the performance of overall framework.
@@ A.6 Exploration of More Text Variants
Additionally, to control variables and compare Entity Grounding performance across different methods, we conduct more challenging tests for RiVEG. RiVEG variants in the w/o Auxiliary Refined Knowledge in MNER group indicate that the text input to their MNER module consists only of the original text and does not incorporate any external knowledge. In this scenario, the MNER performance of these BERT variants of RiVEG is nearly identical to that of the two baseline methods. However, the GMNER and EEG performance of these variants still significantly outperforms all baseline methods. This comparison highlights the robust performance of RiVEG in Entity Grounding.
@@ A.7 Case Study
We further conduct case study to compare the predictions of H-Index and RiVEG for challenging samples. Figure 3 illustrates two examples where H-Index makes errors, while RiVEG produces accurate predictions. For the test sample on the left, even though H-Index correctly predicts the entity-type pairs, it fails to achieve any effective entity grounding. All predictions made by RiVEG are accurate, even for the challenging entity grounding case of the "Preds" team logo. For the sample on the right, H-Index fails to predict all named entities and does not perform entity grounding correctly. RiVEG completes predictions perfectly, even though the grounded area of "Nike SNKRS" is very small.
Figure 4 illustrates two error examples. The left test sample poses a significant challenge for entity grounding due to the resemblance between the shape of the clothing in the image and that of humans. H-Index fails to predict the named entitiy. While RiVEG accurately predicts the entity-type pairs, it incorrectly performs entity grounding. The entity prediction and grounding for the sample on the right are both highly challenging. In this case, H-Index is almost incapable of making effective predictions. RiVEG accurately predicts all entity-type pairs and successfully performs the grounding of "taylor swift". However, it fails to accurately complete the grounding of "the black eyed peas". This indicates that the GMNER task remains highly challenging.
Limitations
In the past, the three-stage reformulation of GMNER was not promising by researchers. Because there may be two intuitive limitations to this reformulation. Limitation 1. Splitting the entire pipeline into three stages can lead to severe error propagation. Limitation 2. Pipeline design seems to be quite heavy. But actually, regarding Limitation 1, this paper demonstrates for the first time that the three-stage pipeline design performs significantly better than existing end-to-end methods. Regarding Limitation 2, the modular nature of RiVEG allows it to freely select lighter sub-models to form more efficient variants to deal with time-critical scenarios. The most lightweight variant of RiVEG can complete the entire three-stage training in 12 hours using a single 4090 GPU, making the training cost quite low. And the inference speed of RiVEG is also acceptable, since LLMs can be called in the form of API. Considering the substantial performance improvement, sacrificing a certain degree of inference speed is worthwhile. In the future, our research direction is to strive for optimal performance while designing each module (VG+VE) in a lightweight style and training them in conjunction, possibly incorporating intermediate auxiliary losses.
Ethics Statement

### v4

@@ 5 Conclusion
In this paper, we propose RiVEG, a unified framework that aims to use LLMs as bridges to reformulate the GMNER task into a unified MNER-VE-VG task. This reformulation enables the framework to seamlessly incorporate the capabilities of state-of-the-art methods for each subtask in their corresponding modules. Extensive experiments show that RiVEG significantly outperforms state-of-the-art methods. The proposed RiVEG greatly reduces the constraints of VG and VE methods on natural language inputs, expands their applicability to more realistic scenarios, and provides insights for future work to better deploy VG and VE modules in related tasks.
Limitations
In the past, the three-stage reformulation of GMNER was not promising by researchers. Because there may be two intuitive limitations to this reformulation. Limitation 1. Splitting the entire pipeline into three stages can lead to severe error propagation. Limitation 2. Pipeline design seems to be quite heavy. But actually, regarding Limitation 1, this paper demonstrates for the first time that the three-stage pipeline design performs significantly better than existing end-to-end methods. Regarding Limitation 2, the modular nature of RiVEG allows it to freely select lighter sub-models to form more efficient variants to deal with time-critical scenarios. The most lightweight variant of RiVEG can complete the entire three-stage training in 5 hours using a single 4090 GPU, making the training cost quite low. And the inference speed of RiVEG is also acceptable, since LLMs can be called in the form of API. Considering the substantial performance improvement, sacrificing a certain degree of inference speed is worthwhile. In the future, our research direction is to strive for optimal performance while designing each module in a lightweight style and training them in conjunction, possibly incorporating intermediate auxiliary losses.
Ethics Statement
Acknowledgments
The authors would like to thank four anonymous reviewers for their insightful comments. This work was supported by the Natural Science Foundation of Tianjin (No.21JCYBJC00640) and the 2023 CCF-Baidu SongGuo Foundation (Research on Scene Text Recognition Based on PaddlePaddle).
@@ A.1 Model Configuration
A.2 Details of Different Module Datasets
| #Twitter-GMNER | #Twitter-GMNER† | # | # | # | # | |
| Train | 7000 | 35000 | 11782 | 58910 | 4694 | 23470 |
| Dev | 1500 | 1500 | 2453 | 2453 | 986 | 986 |
| Test | 1500 | 1500 | 2543 | 2543 | 1036 | 1036 |
| Total | 10000 | 38000 | 16778 | 63906 | 6716 | 25492 |
@@ A.2 Details of Different Module Datasets
Table 7 counts the training data used by different modules. And Table 8 shows the details of the Twitter-GMNER dataset. For the MNER module of RiVEG, the training data for base version comprises original samples and corresponding auxiliary refined knowledge generated by gpt-3.5-turbo provided by PGIM (Li et al., 2023). Additionally, data augmentation is applied to the training set of base version. In data augmentation version, we apply four additional LLMs (i.e., vicuna-7b-v1.5, vicuna-13b-v1.5, llama-2-7b-chat-hf, llama-2-13b-chat-hf) to training set samples using the same algorithm, generating multiple versions of auxiliary refined knowledge for the same sample. The dev set and test set remain the same as base version to ensure the benchmarking of evaluation.
@@ A.5 Exploration of More Visual Variants
We also present the corresponding test set results for different methods of different modules after fine-tuning. The results indicate that stronger combinations of sub-methods generally yield superior performance. The performance of the early classic ALBEF and SeqTR is inferior to the subsequent advanced OFA. Additionally, while the test set Prec@0.5 of SeqTR is slightly lower than that of OFA, there are significant differences in their final GMNER and EEG results. This phenomenon may be attributed to the text representation method employed by SeqTR. The GRU (Chung et al., 2014) vocabulary of SeqTR is solely statistically obtained from or , rendering it incapable of handling out-of-vocabulary words in new samples during inference.
Moreover, data augmentation methods exhibit varying effects on distinct modules. MNER and VG methods can intuitively benefit from data augmentation, but this is not the case for VE methods. One potential reason is that the distribution after fitting more training data does not align with the optimal distribution of the test data. But this is acceptable since data augmentation significantly enhances the performance of overall framework.
@@ A.6 Exploration of More Text Variants
Additionally, to control variables and compare entity grounding performance across different methods, we conduct more challenging tests for RiVEG. The variants in the w/o Auxiliary Refined Knowledge in MNER group indicate that the text input to their MNER module consists only of the original text and does not incorporate any external knowledge. In this scenario, the MNER performance of these BERT variants of RiVEG is nearly identical to that of the two baseline methods. However, the GMNER and EEG performance of these variants still significantly outperforms all baseline methods. This comparison highlights the robust performance of RiVEG in entity grounding.
@@ A.7 Case Study
We further conduct case study to compare the predictions of H-Index (Yu et al., 2023) and RiVEG for challenging samples. Figure 3 illustrates two examples where H-Index makes errors, while RiVEG produces accurate predictions. For the test sample on the left, even though H-Index correctly predicts the entity-type pairs, it fails to achieve any effective entity grounding. All predictions made by RiVEG are accurate, even for the challenging entity grounding case of the “Preds” team logo. For the sample on the right, H-Index fails to predict all named entities and does not perform entity grounding correctly. RiVEG completes predictions perfectly, even though the grounded area of “Nike SNKRS” is very small.
Figure 4 illustrates two error examples. The left test sample poses a significant challenge for entity grounding due to the resemblance between the shape of the clothing in the image and that of humans. H-Index fails to predict the named entity. While RiVEG accurately predicts the entity-type pairs, it incorrectly performs entity grounding. The entity prediction and grounding for the sample on the right are both highly challenging. In this case, H-Index is almost incapable of making effective predictions. RiVEG accurately predicts all entity-type pairs and successfully performs the grounding of “taylor swift”. However, it fails to accurately complete the grounding of “the black eyed peas”. This indicates that the GMNER task remains highly challenging.